[Brahms-l] Shift report 20040311 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu Mar 11 2004 - 07:49:01 EST
------------------------------------------------------------
Date: Thu Mar 11 07:49:00 2004, Shift: 00:00-08:00
Supervisor: CEJ, with: FH
Log book#15 pages 83 - 83

Shift Summary:
------------------------------------------------------------
1) Beam from 3:20 to 6:20 and again from 7:30 (beam studies)

2) Data runs:
   10898; tr2: 41K, tr3:  46K
   10899; tr2: 50K, tr3:  75K
   10902; tr2: 40K, tr3: 100K 
   10903; tr2: 18K, tr3:  46K
   10904; still running (beam studies)

3) No problems reported (T3-2-X1 dist still looks bad).
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 11 07:49:10 2004

This archive was generated by hypermail 2.1.8 : Thu Mar 11 2004 - 07:49:32 EST